# Truth Knows No Language: Evaluating Truthfulness Beyond English

### v2

@@ 6 Related Work
A significant challenge in contemporary Artificial Intelligence (AI) research concerns the development of methodologies to optimize LLMs for factual accuracy and veracity in their outputs. Improving factual consistency and reducing hallucinations would help to increase trust in LLMs thereby increasing their application across various domains. Apart from the popular TruthfulQA, already introduced in Section 2, other approaches include SimpleQA Wei et al. (2024), and VeritasQA Aula-Blasco et al. (2025).
SimpleQA is a benchmark dataset designed for evaluating the abilities of LLMs to answer factual questions, specifically targeting short, fact-seeking queries. The dataset features dual-source verification for answer validation, and shows an increased difficulty compared to legacy benchmarks (e.g., Joshi et al. (2017) or Kwiatkowski et al. (2019)), where current LLMs show performance saturation.
@@ 7 Conclusion
Our expansion to include Iberian languages certainly only covers a small fraction of the world’s languages, limiting the generalizability of our findings. However, our results using automatic translation suggest potential scalability to other languages, offering a cost-effective approach to broadening language coverage of the benchmark without the need to get professional translations.
The challenge of evaluating generative tasks automatically is another significant limitation. On the one hand, a fully manual evaluation was deemed impractical due to the significant burden it would impose on resources and scalability. On the other hand, although automated methods like LLM-as-judge work well, they still fall short of capturing the accuracy and subtle nuances that manual evaluations can achieve — especially as models continue to improve.
@@ Acknowledgments
We are thankful to the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and by FEDER, EU; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iii) NÓS-ILENIA (2022/TL22/0021533). Also, Xunta de Galicia: Centro de investigación de Galicia accreditation 2024-2027 ED431G-2023/04. Blanca Calvo Figueras is supported by the UPV/EHU PIF22/84 predoc grant. Julen Etxaniz holds a PhD grant from the Basque Government (PRE_2024_2_0028). Iria de Dios Flores is supported by a Juan de la Cierva contract and project (JDC2022-049433-I, financed by the MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR”.
@@ Appendix C Translation Metrics
Table 7: Evaluation of the machine-translated TruthfulQA dataset using the human translations as reference.
@@ Appendix E Comparing the results of two Judge-LLMs from different families
Table 8: Ranking of the TruthfulQA results with two Judge-LLMs of different families: Llama-3.1 and Gemma-2. They show that family of the judge does not influence the results.

### v3

@@ 6 Related Work
A significant challenge in contemporary Artificial Intelligence research concerns the development of methodologies to optimize LLMs for factual accuracy and veracity in their outputs. Improving factual consistency and reducing hallucinations would help to increase trust in LLMs, thereby increasing their application across various domains. Apart from the popular TruthfulQA, already introduced in Section 2, other approaches include SimpleQA Wei et al. (2024), and VeritasQA Aula-Blasco et al. (2025).
SimpleQA is a benchmark dataset designed for evaluating the abilities of LLMs to answer factual questions, specifically targeting short, fact-seeking queries. The dataset features dual-source verification for answer validation and shows an increased difficulty compared to legacy benchmarks (e.g., Joshi et al. (2017) or Kwiatkowski et al. (2019)), where current LLMs show performance saturation.
@@
Finally, prior research has demonstrated that multilingual models exhibit inconsistencies in factual knowledge across different languages Wang et al. (2024); Liu et al. (2025), and that these inconsistencies persist regardless of model size Qi et al. (2023). Moreover, instruction-tuning LLMs in multilingual contexts remains an underexplored area Lai et al. (2023), contributing to the uneven performance of current multilingual LLMs.
@@ 7 Conclusion
Our expansion to include Iberian languages certainly only covers a small fraction of the world’s languages, limiting the generalizability of our findings. However, our results using automatic translation suggest potential scalability to other languages, offering a cost-effective approach to broadening the language coverage of the benchmark without the need to get professional translations.
The challenge of evaluating generative tasks automatically is another significant limitation. On the one hand, a fully manual evaluation was deemed impractical due to the significant burden it would impose on resources and scalability. On the other hand, although automated methods like LLM-as-judge work well, they still fall short of capturing the accuracy and subtle nuances that manual evaluations can achieve, especially as models continue to improve.
@@ Acknowledgments
We are thankful to the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and by FEDER, EU; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iv) NÓS-ILENIA (2022/TL22/0021533) and LingUMT, grant PID2021-128811OA-I00, MEC. Also, Xunta de Galicia: Centro de investigación de Galicia accreditation 2024-2027 ED431G-2023/04. Blanca Calvo Figueras is supported by the UPV/EHU PIF22/84 predoc grant. Julen Etxaniz holds a PhD grant from the Basque Government (PRE_2024_2_0028). Iria de-Dios-Flores is supported by the project JDC2022-049433-I, financed by the MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR”, and grant SGR 2021 00470, financed by AGAUR (Catalan Government).
@@ Appendix C Translation Metrics
Table 7: Evaluation of the machine-translated TruthfulQA dataset using the human translations as reference. All the metrics were computed using the Hugging Face implementation in https://huggingface.co/docs/evaluate.
@@ Appendix E Comparing the results of two Judge-LLMs from different families
Table 8: Ranking of the TruthfulQA results with two Judge-LLMs of different families: Llama 3.1 and Gemma 2. They show that family of the judge does not influence the results.